if you invest 15,000 in an account paying 4% compounded annually what will the balance be after 9 years
Answer by lwsshak3(11628)

if you invest 15,000 in an account paying 4% compounded annually what will the balance be after 9 years
Compound Interest formula: A=P(1+i)^n, P=initial investment, i=interest rate per period, n=number of periods, A=amount after n-periods.
For given problem:
Initial investment=15000
interest rate per period=4%
number of periods=9
..
A=15000(1+.04)^9=15000(1.04^9)=21350
Balance after 9 years=21,350